OpenStudy (anonymous):

(x^2-xy)/(y-x). I know the answer is -x I just don't understand why.

OpenStudy (phi):

did you factor x from the top ? and factor -1 from the bottom

OpenStudy (amistre64):

\[\frac{x^2-xy}{y-x}\] \[\frac{x^2-xy+y^2-y^2-xy+xy}{y-x}\] \[\frac{(x-y)(x-y)-y^2+xy}{y-x}\] \[\frac{(y-x)(x-y)+y^2-xy}{y-x}\] \[(x-y)+\frac{y^2-xy}{y-x}\] \[(x-y)+\frac{y(y-x)}{y-x}\] \[(x-y)+y\] review that for mistakes of course

OpenStudy (amistre64):

phis idea might be simpler ... but not as spectacular ;)

OpenStudy (phi):

spectacular is not the adjective that springs to mind...

OpenStudy (phi):

but it does work...

OpenStudy (phi):

but there is a minus sign missing...

OpenStudy (amistre64):

in the 4th step i pulled it out of the square and for some reason thought it distrbuted over the -y^2 + xy ... in error

OpenStudy (anonymous):

@amistre64 Although spectacular looking, I still do not understand your explanation... :c and @phi i did what you said but I don't see how it gets me -x. D:

OpenStudy (phi):

jayycee post your work

OpenStudy (amistre64):

my method amounts to completing a square is all; which you may not have covered yet

OpenStudy (anonymous):

(x(x-xy))/(-1(-y+x))

OpenStudy (phi):

if we distribute x(x-xy) with get x^2 -x^2 y which not what you started with redo the top

OpenStudy (anonymous):

x(x-y) / -1 (-y+x) ?

OpenStudy (phi):

yes, now x(x-y) = x^2 - xy which matches. notice you can write -y+x as x+ -y or x-y can you simplify what you have...

OpenStudy (anonymous):

Sooo, x(x-y) / _y +x ? but then what?

OpenStudy (phi):

? you start with \[ \frac{x(x-y)}{-1(-y+x) } \] or , writing -y+x as x-y \[ \frac{x(x-y)}{-1(x-y) } \] remember anything divided by itself is 1.

OpenStudy (anonymous):

OH! how did I forget!. Thank you so much!

OpenStudy (phi):

do you know how to multiply fractions? top times top and bottom times bottom that means that \[ \frac{x}{-1} \cdot \frac{(x-y)}{(x-y)} = \frac{x(x-y)}{-1(x-y)} \] but what is \[ \frac{(x-y)}{(x-y)} \] ?

OpenStudy (anonymous):

just one

OpenStudy (phi):

so you are left with x/-1 which is -x (one way is multiply top and bottom by -1 to get -x/1 or just -x)

OpenStudy (phi):

is something unclear ?

OpenStudy (anonymous):

Nope! i completely understand!:D